OpenStudy (fatboi96):

What is the rate of change and initial value for the linear relation that includes the points shown in the table? x y 1 20 3 10 5 0 7 -10 Initial value: 20, rate of change: 10 Initial value: 30, rate of change: 10 Initial value: 25, rate of change: -5 Initial value: 20, rate of change: -10

OpenStudy (jiteshmeghwal9):

Let's find out the linear relation let the linear relation be y=mx+c . using the given information at x=1 & y=20, 20=m+c .....(1) at x=5 & y=0 , 0=5m+c......(2) From the equation (1) & (2) m=-5 , c=25 so the linear relation comes out to be y=-5x+25

OpenStudy (jiteshmeghwal9):

at x=0 u will gt the initial value of y & the rate of change is the slope

OpenStudy (fatboi96):

so its c